# Domain Generalizable Knowledge Tracing via Concept Aggregation and Relation-Based Attention

### v1

@@ III-E Generalization to Target Domain
Figure 4: A toy example of relation-based attention, where different shapes represent different concepts and different color represent different exercises. Relation-based attention focus more on related exercises and concepts.
IV Domain-Generalizable Relation-based Knowledge Tracing
Although our proposed DGKT framework effectively enhances the performance of existing KT models in new domains, these models are not inherently designed for domain generalization in knowledge tracing. One of the biggest challenges hindering these models from transferring to a new domain is the complete discrepancy in exercise IDs across domains, which leads to an inability to fully leverage the information from exercise IDs in students’ interaction sequences. To address this issue, we propose a domain-generalizable relation-based knowledge tracing model(DGRKT) specifically designed for domain generalization in knowledge tracing. This model captures the relationships between exercises and concepts across different timesteps based on attention mechanism. In this section, we illustrate DGRKT in detail including DGRKT model architecture and relation-based attention encoder.
IV-A DGRKT Model Architecture
Similar to the aforementioned model architecture, our DGRKT consists of a feature embedding module, a knowledge state encoder, and a knowledge state decoder. In DGRKT, the feature embedding module and knowledge state decoder remain the same, while the knowledge state encoder is specifically designed to capture the relationships between different timesteps. As described above, we can transform a student’s interaction sequence into question embedding matrix and question-response embedding matrix:
$\begin{split}M_{q},M_{qr}=FeatureEmbeding((q_{1},r_{1}),(q_{2},r_{2}),...,(q_{t},r_{t})).\\ \end{split}$ (27)
@@ IV-A DGRKT Model Architecture
Given a question embedding matrix $M_{q}$ and a question-response embedding matrix $M_{qr}$, the outputted knowledge state can be represented as:
$\begin{split}x_{t}&=\text{R-Attn}(\mathcal{Q}=M_{q},\mathcal{K}=M_{q},\mathcal{V}=M_{qr}),\\ h_{t}&=\text{SeqIN}(x_{t}),\end{split}$ (28)
@@
Likewise, the probability of a student answering the next question correctly can be calculated as:
$\begin{split}\hat{y}_{t+1}&=\text{Dec}(h_{t+1},q_{t+1}).\end{split}$ (29)
Same as before, we conduct concept aggregation to realize generalization to the target domain.
IV-B Relation-based Attention Encoder
To explicitly capture the relation between different timesteps and fully leverage exercise IDs from different domains, we have designed a novel relation-based attention encoder that adjusts the attention value $\alpha$ based on the relation between two timesteps. We assume that the more similar a question a student answered at a previous time is to the current question, the more valuable the information from that previous time will be. Therefore, we have defined three types of relations based on the relevance between two timesteps to guide the learning of the relation-based attention encoder, as is shown in Fig. 4.
Specifically, we define three kinds of relations between two timesteps. Given two timesteps $i$ and $j$, and their corresponding interactions $(q_{i},r_{i})$ and $(q_{j},r_{j})$, the relevance between two exercises $q_{i}$ and $q_{j}$ can be represented as:
where $R(i,j)$ denotes the relation between timestep $i$ and timestep $j$. $R(i,j)=0$ indicates $q_{i}$ and $q_{j}$ are irrelevant, $R(i,j)=1$ indicates $q_{i}$ and $q_{j}$ have common related concepts, and $R(i,j)=2$ indicates $q_{i}$ and $q_{j}$ are the same. $C_{i}$ and $C_{j}$ are the concept sets associated with $q_{i}$ and $q_{j}$, respectively, consisting of related concepts.
Following the procedure of attention mechanism, we assume that $e_{q}r$ corresponds to value and $e_{q}$ corresponds to query and key. In attention block, the attention value $\alpha_{ij}$ is calculated as:
$\begin{split}q_{i}&=W_{q}e_{q_{i}},\\ k_{j}&=W_{k}e_{q_{j}},\\ \alpha_{ij}&=Softmax(\frac{q_{i}^{T}k_{j}}{\sqrt{d}}),\end{split}$ (31)
where $W_{q}$ and $W_{k}\in\mathbb{R}^{d\times d}$ are linear transformation, $d$ is the dimension of features.
Normally, in an attention mechanism, the calculated attention value $\alpha$ is directly used to compute the weighted sum. However, in the relation-based attention encoder, our model aims to explicitly differentiate between different relations. Therefore, we adjust the attention value $\alpha_{ij}$ based on the relation $R(i,j)$. Specifically, for timestep $j$, its attention values $[\alpha_{1j},\alpha_{2j},...,\alpha_{(j-1)j}]$ are modified as follows:
$\begin{split}v_{i}&=W_{v}e_{q_{i}},\\ \lambda_{ij}&=\begin{cases}a&\text{if }R(i,j)=0\\ b&\text{if }R(i,j)=1\\ c&\text{if }R(i,j)=2\end{cases},\\ A&=[\lambda_{1j}\alpha_{1j},\lambda_{2j}\alpha_{2j},...,\lambda_{(j-1)j}\alpha_{(j-1)j}],\\ x_{j}&=\sum_{i=1}^{j-1}\frac{\lambda_{ij}\alpha_{ij}}{\sum A}v_{i},\end{split}$ (32)
where $W_{v}\in\mathbb{R}^{d\times d}$ is a linear transformation for the value vectors $v_{i}$. $\lambda_{ij}$ represents the importance of timestep $i$ for timestep $j$. It is noted that $0<a<b<c$ to ensure that $\lambda_{ij}$ increases with the relevance of $i$ and $j$. The output of the relation-based attention encoder, denoted as $x$, is computed as the weighted sum of the value vectors $v$.
V Experiments
In this section, we evaluate the performance of our proposed method using five well-known knowledge tracing benchmark datasets on domain generalization knowledge tracing. Also, we conduct visualization experiment along with cluster analysis that manifest the effectiveness of our proposed concept aggregation.
TABLE I: The AUC results of four different knowledge tracing methods on five target domain datasets with only a few interactions in the target domain and sufficient interaction sequences in the other four source domains.
Dataset batch num DKT DG-DKT SAINT DG-SAINT AKT DG-AKT SimpleKT DG-SimpleKT DGRKT
ASSIST09 1 0.683 0.721 0.706 0.718 0.704 0.718 0.706 0.713 0.729
2 0.685 0.727 0.711 0.727 0.713 0.723 0.713 0.715 0.737
4 0.687 0.732 0.716 0.731 0.716 0.729 0.716 0.720 0.740
8 0.701 0.734 0.719 0.735 0.720 0.736 0.720 0.721 0.743
ASSIST15 1 0.658 0.694 0.680 0.683 0.670 0.690 0.670 0.691 0.717
2 0.671 0.705 0.691 0.699 0.684 0.701 0.684 0.692 0.729
4 0.675 0.710 0.696 0.708 0.691 0.705 0.691 0.695 0.732
8 0.677 0.716 0.702 0.714 0.694 0.719 0.694 0.696 0.732
ASSIST17 1 0.600 0.675 0.636 0.649 0.636 0.661 0.636 0.655 0.677
2 0.607 0.681 0.647 0.654 0.650 0.673 0.650 0.664 0.684
4 0.614 0.685 0.655 0.661 0.661 0.683 0.661 0.669 0.689
8 0.617 0.693 0.661 0.667 0.669 0.685 0.669 0.674 0.696
ALGEBRA 1 0.693 0.749 0.751 0.765 0.741 0.756 0.741 0.748 0.775
2 0.709 0.760 0.756 0.772 0.752 0.775 0.752 0.757 0.781
4 0.722 0.764 0.762 0.779 0.761 0.781 0.761 0.769 0.784
8 0.732 0.776 0.765 0.785 0.764 0.790 0.764 0.775 0.795
Junyi 1 0.607 0.678 0.688 0.699 0.678 0.679 0.678 0.685 0.724
2 0.637 0.684 0.689 0.701 0.686 0.692 0.686 0.688 0.727
4 0.659 0.689 0.692 0.708 0.691 0.702 0.691 0.699 0.728
8 0.662 0.691 0.692 0.711 0.695 0.709 0.695 0.702 0.730
V-A Datasets and setting.
We evaluate the performance of our DGKT using five well-known KT benchmark datasets: ASSISTment 2009111https://sites.google.com/site/assistmentsdata/home/assistment-2009-2010-data/skill-builder-data-2009-2010, ASSISTment 2015222https://sites.google.com/site/assistmentsdata/home/2015-assistments-skill-builder-data, ASSISTment 2017333https://sites.google.com/view/assistmentsdatamining/dataset, Algebra 2005444https://pslcdatashop.web.cmu.edu/KDDCup/, and Junyi [^5] http://www.junyiacademy.org/. The first three datasets are provided by an online education system called ASSISTment [38] with different exercises and students, which are widely used to evaluate KT models. Junyi [39] was collected from an E-learning platform called Junyi Academy, which was established in 2012. Algebra 2005 is provided by the KDDcup 2010 Educational Data Mining challenge containing 13–14 year old students’ interaction on Algebra questions. For convenience, these datasets are abbreviated as ASSIST09, ASSIST15, ASSIST17, STATICS and ALGEBRA in the following.
In the DGKT model, after conducting a grid search, the dimension $d$ of feature embeddings is determined to be 256, and the number $k$ of concept clusters is set to 5. The optimization of the model employs the Adam optimizer with a learning rate of 0.0001, and the batch size is fixed at 32. In relation-based attention, a, b, c are set to 1, 2 and 4 respectively.
Our approach adopts a domain generalization setup, where one dataset serves as the target domain, while the remaining four datasets act as four source domains. For concept feature learning, we conduct 12,000 epochs of training on the source domains. In each epoch, every source domain is sampled for a batch, and the model is trained on four batches of data. Subsequently, for concept centroid refinement, we similarly conduct 6,000 epochs of training.
On the target domain, we test our method with different scale of training data, from a single batch of data to 8 batches of data, and a fine-tuning process of 15 epochs is conducted. This approach allows us to assess the effectiveness of knowledge tracing across all five datasets.
V-B Comparison Results and Analysis
We present a comparative analysis between our proposed DGKT method and four prominent knowledge tracing models, i.e., DKT [6], SAINT [8], AKT [9], and SimpleKT [40]. For fair comparison, we initially pre-train these KT models on all available source domains, followed by fine-tuning them on the specific target domain. Methods including DKT, AKT, and SimpleKT are reimplemented by their public source codes, while SAINT is reimplemented based on their papers. Also, the hyperparameters are set as the values provided by the literature.
Domain Generalization of Knowledge Tracing Tasks. The AUC results for all compared KT methods in the context of domain generalization are presented in Table I. DKT, SAINT, AKT and SimpleKT are pretrained on all source domains and finetuned on target domain, while DG-DKT, DG-SAINT, DG-AKT and DG-SimpleKT are corresponding KT models with DGKT framework. These results highlight the effectiveness of the DGKT approach in enhancing knowledge tracing performance with limited training data. In the target domain, DGKT demonstrates significant improvements due to its robust generalization capabilities. Addidtionally, we find that DG-DKT performs better than other models based DGKT framework in several dataset despite its relatively small parameter size. We believe it is possibly because its relatively small parameter size makes DG-DKT better generalization abilility. All knowledge tracing models integrated with the DGKT framework exhibit superior performance compared to traditional knowledge tracing models pretrained on source domains. As the size of the data increases, our method consistently maintains its advantage.
TABLE II: The AUC results of n different knowledge tracing methods on five target domain datasets with only a few interactions in the target domain and sufficient interaction sequences in the other four source domains.
Dataset batch num DG-DKT -w/o SeqIN DG-DKT -BN DG-DKT -LN DG-DKT -SeqIN
ASSIST09 1 0.712 0.718 0.713 0.721
2 0.717 0.722 0.725 0.727
4 0.726 0.730 0.727 0.732
8 0.731 0.729 0.730 0.734
ASSIST15 1 0.689 0.691 0.688 0.694
2 0.700 0.702 0.704 0.705
4 0.704 0.710 0.713 0.710
8 0.706 0.715 0.719 0.716
ASSIST17 1 0.657 0.671 0.651 0.679
2 0.677 0.679 0.671 0.689
4 0.680 0.691 0.679 0.695
8 0.690 0.697 0.695 0.703
ALGEBRA 1 0.715 0.748 0.746 0.749
2 0.743 0.759 0.752 0.760
4 0.764 0.762 0.760 0.764
8 0.765 0.775 0.772 0.776
Junyi 1 0.634 0.668 0.659 0.678
2 0.653 0.669 0.679 0.684
4 0.681 0.680 0.677 0.689
8 0.682 0.675 0.679 0.691
$\mathcal{A}-distance=1.82$
@@ V-B Comparison Results and Analysis
Figure 5: The $t$-SNE visualization of feature embedding of student interactions by using different normalization methods.
V-C Ablation Studies.
Concept Aggregation. We conducted ablation experiments to investigate the impact of our proposed concept aggregation methods, which include concept feature learning (CFL), centroid refinement (CR), and target embedding adaptation (TA). Specifically, we compare our DGKT model with three variants: DGKT-w/o CFL, DGKT-w/o CR, and DGKT-w/o TA. The details of these variants are presented below:
  • DGKT-w/o CFL: This variant omits concept feature learning. The model is directly trained on the target domain without pretraining on the source domain.
  • DGKT-w/o CR: This variant omits centroid refinement. After concept feature learning, the model is directly fine-tuned on the target domain without centroid refinement.
  • DGKT-w/o TA: This variant omits target embedding adaptation. The model is fine-tuned on the target domain using randomly initialized embeddings, and question embeddings are calculated solely by Eqs. (25).
From Table III, we can observe that DGKT without Concept Feature Learning (CFL) shows a significant decrease in AUC. This indicates that CFL is a crucial procedure, as it derives meta-knowledge from several source domains. Additionally, concept aggregation proves to be beneficial for learning a robust concept embedding for DGRKT, as well as for adapting the target embedding effectively.
Sequence Instance Normalization. Moreover, we conduct ablation experiment on Sequence Instance Normalization (SeqIN), we compare DG-DKT with three variants: DG-DKT -w/o SeqIN (SeqIN removed), DG-DKT -BN (SeqIN replaced by BN) and DG-DKT -LN (SeqIN replaced by LN).
We omit instance normalization because its calculation inevitably includes all timesteps, leading to information leakage in knowledge tracing. The results, shown in Table II, demonstrate the advantage of SeqIN. Specifically, DG-DKT without SeqIN (DG-DKT -w/o SeqIN) shows a significant decrease in AUC, indicating that the normalization method is crucial in such cross-domain tasks and Sequence Instance Normalization is a better normalization module for domain generalization of knowledge tracing.
TABLE III: The AUC results of n different knowledge tracing methods on five target domain datasets with only a few interactions in the target domain and sufficient interaction sequences in the other four source domains.
Dataset batch num DGRKT -w/o CFL DGRKT -w/o CR DGRKT -w/o TA DGRKT
ASSIST09 1 0.699 0.711 0.709 0.729
2 0.705 0.717 0.718 0.737
4 0.716 0.725 0.724 0.740
8 0.726 0.728 0.729 0.743
ASSIST15 1 0.670 0.697 0.695 0.717
2 0.689 0.709 0.712 0.729
4 0.707 0.718 0.717 0.732
8 0.719 0.723 0.723 0.732
ASSIST17 1 0.624 0.646 0.645 0.677
2 0.649 0.663 0.662 0.684
4 0.667 0.672 0.672 0.689
8 0.675 0.681 0.680 0.696
ALGEBRA 1 0.701 0.750 0.752 0.775
2 0.756 0.762 0.763 0.781
4 0.764 0.767 0.767 0.784
8 0.772 0.774 0.780 0.795
Junyi 1 0.691 0.700 0.697 0.724
2 0.704 0.706 0.707 0.727
4 0.709 0.712 0.711 0.728
8 0.713 0.714 0.713 0.730
V-D Visualization
Visualization of SeqIN. To further demonstrate the effectiveness of our SeqIN, we visualize the knowledge state $h$ via $t$-SNE [41] with different normalization methods in Fig. 5. This figure shows four different source domains (ASSIT15, ASSIT17, Junyi and Algebra05), and each color represents one source domain. Moreover, we use $Proxy\mathcal{A}-distance$ to calculate the divergence between different domains [42].
From Fig. 5 we can observe that the SeqIN can significantly aggregate the feature embeddings from different source domains, resulting in the lowest $\mathcal{A}-distance$, while the BN and LN do not work well on the embedding sequences of student interactions. This shows the efficacy of our SeqIN again to reduce the domain discrepancy for knowledge tracing.
Figure 6: A visualization of student’s mastery on five concept clusters for a student’s interaction randomly chosen from ASSIST09.
TABLE IV: Visualization of concept aggregation results.
@@ V-D Visualization
4 Venn Diagram, Counting Methods, Congruence, Proportion, Prime Number, Algebraic , implification, Equation Solving More Than Two Steps, Write Linear Equation from Situation

### v2

@@ 3.4 Relation-Aware Knowledge Encoder
Figure 5: The architecture of relation-aware attention. Relation-aware attention focus more on related exercises and concepts.
Given two timesteps $i$ and $j$, and their corresponding interactions $(q_{i},r_{i})$ and $(q_{j},r_{j})$, the relevance $R(i,j)$ between two questions $q_{i}$ and $q_{j}$ can be represented as:
where $C_{i}$ and $C_{j}$ are the concept sets associated with questions $q_{i}$ and $q_{j}$, respectively, and $C_{i}\cap C_{j}\neq\emptyset$ implies the presence of shared concepts. Thus, $R(i,j)=0$ indicates that $q_{i}$ and $q_{j}$ are irrelevant, $R(i,j)=1$ indicates that the two questions may be relevant due to their association with shared concepts, and $R(i,j)=2$ indicates $q_{i}$ and $q_{j}$ are the same question.
Following the procedure of attention mechanism, we employ $e_{q}$ to corresponds to query ($Q$) and key ($K$), and $e_{qr}$ corresponds to value ($V$). In the attention block, the attention value $\alpha_{ij}$ is calculated as:
$\begin{split}Q_{i}&=W_{q}e_{q_{i}},\\ K_{j}&=W_{k}e_{q_{j}},\\ \alpha_{ij}&=\text{Softmax}(\frac{Q_{i}^{T}K_{j}}{\sqrt{d}}),\end{split}$ (28)
where $W_{q}$ and $W_{k}\in\mathbb{R}^{d\times d}$ are the linear transformation matrices, and $d$ is the dimension of features.
Typically, within the attention mechanism, the attention value $\alpha$ is directly used to compute the weighted sum. In contrast, the relation-aware attention encoder in our model seeks to explicitly distinguish among various relationships. Consequently, we modify the attention value $\alpha_{ij}$ based on the relevance $R(i,j)$. Specifically, for the timestep $j$, the adjusted attention value $\boldsymbol{\tilde{\alpha}}_{j}$ is calculated as follows:
$\begin{split}\lambda_{ij}&=\begin{cases}a&\text{if }R(i,j)=0\\ b&\text{if }R(i,j)=1\\ c&\text{if }R(i,j)=2\end{cases},\\ \boldsymbol{\tilde{\alpha}}_{j}&=[\lambda_{1j}\alpha_{1j},\lambda_{2j}\alpha_{2j},...,\lambda_{(j-1)j}\alpha_{(j-1)j}],\\ \end{split}$ (29)
where $\lambda_{ij}$ signifies the relevance weight, and $a,b$ and $c$ is the learned parameters. Let $0<a<b<c$, this condition ensures that $\lambda_{ij}$ increases with the relevance $R(i,j)$. Subsequently, the weighted sum can be computed using the adjusted attention value as follows:
$\begin{split}V_{i}&=W_{v}e_{qr_{i}},\\ x_{j}&=\sum_{i=1}^{j-1}\frac{\lambda_{ij}\alpha_{ij}}{\sum\boldsymbol{\tilde{\alpha}}_{j}}V_{i},\end{split}$ (30)
where $W_{v}\in\mathbb{R}^{d\times d}$ is the linear matrix for the value vectors $V_{i}$, and $x_{j}$ represents the output of the relation-aware attention encoder for the timestep $j$.
By replacing the knowledge state encoder with the RA-Encoder in the DGKT framework, we introduce the Domain-Generalizable Relation-Aware Knowledge Tracing model (DGrKT).
@@
Given a question embedding matrix $\textbf{M}_{q}$ and a question-response embedding matrix $\textbf{M}_{qr}$, the outputted knowledge state can be represented as:
$\begin{split}x_{t}&=\text{RA-Encoder}(\mathcal{Q}=\textbf{M}_{q},\mathcal{K}=\textbf{M}_{q},\mathcal{V}=\textbf{M}_{qr}),\\ h_{t}&=\text{SeqIN}(x_{t}),\end{split}$ (31)
where $\mathcal{Q}$, $\mathcal{K}$, and $\mathcal{V}$ represent the query, key, and value, respectively, in the attention mechanism. $\text{SeqIN}(\cdot)$ is the aforementioned sequence instance normalization. As discussed earlier, the knowledge state decoder can then transform $h_{t}$ into the probability of a student answering the next question correctly. Consistent with our DGKT, we apply concept aggregation to facilitate generalization to the target domain.
4 Experiments
In this section, we evaluate the performance of our proposed DGKT framework and DGrKT model using five well-known knowledge tracing benchmark datasets on multi-source domain adaptation knowledge tracing. Also, we conduct visualization experiments along with cluster analysis that manifest the effectiveness of our proposed concept aggregation.
4.1 Datasets and Setting
We evaluate the performance of our DGKT framework and DGrKT using five well-known KT benchmark datasets: ASSISTment 2009111https://sites.google.com/site/assistmentsdata/home/assistment-2009-2010-data/skill-builder-data-2009-2010, ASSISTment 2015222https://sites.google.com/site/assistmentsdata/home/2015-assistments-skill-builder-data, ASSISTment 2017333https://sites.google.com/view/assistmentsdatamining/dataset, Algebra 2005444https://pslcdatashop.web.cmu.edu/KDDCup/, and Junyi [^5] http://www.junyiacademy.org/. The first three datasets are provided by an online education system called ASSISTment [14] with different exercises and students, which are widely used to evaluate KT models. Junyi [8] was collected from an E-learning platform called Junyi Academy, which was established in 2012. Algebra 2005 is provided by the KDDcup 2010 Educational Data Mining challenge containing 13–14 year old students’ interaction on Algebra questions. Each dataset is composed of extensive student interaction records collected from its respective educational platform. Based on these records, we constructed a personalized sequence of question-response interactions for each student, which were then segmented into fixed-length sequences of 200 interactions. To ensure data quality, students with fewer than 20 total interactions were excluded. The final dataset was partitioned into training and testing sets using an 80:20 split. For convenience, these datasets are abbreviated as ASSIST09, ASSIST15, ASSIST17, ALGEBRA and Junyi in the following.
Table 1 summarizes the statistics of these datasets. The number of students ranges from 565 (ALGEBRA05) to 10,451 (ASSIST15), and the number of unique problems varies significantly, with ALGEBRA05 containing the most (211,397) and ASSIST15 containing the least (100). The total number of student attempts also varies widely, with Junyi having the largest volume (4,371,160 attempts), while ASSIST09 has the least (330,116 attempts). These datasets present different levels of data sparsity and distribution shifts, making them ideal benchmarks for evaluating the cross-domain capability of our proposed method.
In the DGKT framework based model, after conducting a grid search, the dimension $d$ of feature embeddings is determined to be 256, and the number $k$ of concept clusters is set to 5. The optimization of the model employs the Adam optimizer with a learning rate of 0.0001, and the batch size is fixed at 32.
Our approach adopts a multi-source domain adaptation setup, where one dataset serves as the target domain, while the remaining four datasets act as four source domains. For concept feature learning, we conduct 12,000 epochs of training on the source domains. In each epoch, every source domain is sampled for a batch and the model is trained on four batches of data. Subsequently, for concept prototypes refinement, we similarly conduct 6,000 epochs of training.
On the target domain, we test our method with different scales of training data, from a single batch of data to 8 batches of data, and a fine-tuning process of 50 epochs is conducted. This approach allows us to assess the effectiveness of knowledge tracing across all five datasets.
Table 1: Statistics of the five benchmark datasets used in our experiments.
Dataset # Students # Problems # Concepts # Attempts
ASSIST09 2314 26521 124 330116
ASSIST15 10451 100 100 624487
ASSIST17 1705 3162 102 942771
ALGEBRA05 565 211397 113 813569
Junyi 10000 707 40 4371160
4.2 Comparison Results and Analysis
We present a comparative analysis between our proposed DGKT framework, the DGrKT model, and five prominent knowledge tracing (KT) models, i.e., DKT [28], SAINT [10], AKT [15], SimpleKT (abbreviated a s SKT)[25], and RobustKT (abbreviated as RBKT)[16]. DG-DKT, DG-SAINT, DG-AKT, DG-SKT, and DG-RKT are instantiated from our DGKT framework, corresponding to the five KT models above. For fair comparison, we initially pre-train these KT models on all available source domains, followed by fine-tuning them on the specific target domain. Methods including DKT, AKT, SimpleKT and RobustKT are reimplemented by their public source codes, while SAINT is reimplemented based on their papers. Also, the hyperparameters are set as the values provided by the literature.
Table 2: The AUC results of four different knowledge tracing methods on five target domain datasets with only a few interactions in the target domain and sufficient interaction sequences in the other four source domains.
Datasets batch DKT DG-DKT SAINT DG-SAINT AKT DG-AKT SKT DG-SKT RBKT DG-RBKT DGrKT
ASSIST09 1 0.683 0.721 0.706 0.718 0.704 0.718 0.706 0.713 0.695 0.706 0.729
2 0.685 0.727 0.711 0.727 0.713 0.723 0.713 0.715 0.704 0.715 0.737
4 0.687 0.732 0.716 0.731 0.716 0.729 0.716 0.720 0.711 0.720 0.740
8 0.701 0.734 0.719 0.735 0.720 0.736 0.720 0.721 0.719 0.729 0.743
ASSIST15 1 0.658 0.694 0.680 0.683 0.672 0.690 0.670 0.691 0.668 0.675 0.717
2 0.671 0.705 0.691 0.699 0.682 0.701 0.684 0.692 0.674 0.685 0.729
4 0.675 0.710 0.696 0.708 0.687 0.705 0.691 0.695 0.695 0.701 0.732
8 0.677 0.716 0.702 0.714 0.698 0.719 0.694 0.696 0.705 0.717 0.732
ASSIST17 1 0.600 0.675 0.636 0.649 0.640 0.661 0.636 0.655 0.627 0.641 0.677
2 0.607 0.681 0.647 0.654 0.656 0.673 0.650 0.664 0.632 0.660 0.684
4 0.614 0.685 0.655 0.661 0.656 0.683 0.661 0.669 0.661 0.681 0.689
8 0.617 0.693 0.661 0.667 0.664 0.685 0.669 0.674 0.678 0.691 0.696
ALGEBRA 1 0.693 0.749 0.751 0.765 0.739 0.756 0.741 0.748 0.688 0.717 0.775
2 0.709 0.760 0.756 0.772 0.755 0.775 0.752 0.757 0.717 0.749 0.781
4 0.722 0.764 0.762 0.779 0.759 0.781 0.761 0.769 0.768 0.771 0.784
8 0.732 0.776 0.765 0.785 0.765 0.790 0.764 0.775 0.777 0.784 0.795
Junyi 1 0.607 0.678 0.688 0.699 0.673 0.679 0.678 0.685 0.664 0.662 0.724
2 0.637 0.684 0.689 0.701 0.688 0.692 0.686 0.688 0.678 0.685 0.727
4 0.659 0.689 0.692 0.708 0.692 0.702 0.691 0.699 0.687 0.691 0.728
8 0.662 0.691 0.692 0.711 0.697 0.709 0.695 0.702 0.694 0.696 0.730
Multi-source Domain Adaptation for Cold-start Knowledge Tracing Tasks. The AUC results for all compared KT methods in the multi-source domain adaptation are presented in Table 2.These results highlight the effectiveness of the DGKT framework in enhancing knowledge tracing performance with limited training data. In the target domain, DGKT framework demonstrates significant improvements due to its robust generalization capabilities. All knowledge tracing models integrated with the DGKT framework exhibit superior performance compared to traditional knowledge tracing models pretrained on source domains. Moreover, our DGrKT performs the best due to its specially designed attention mechanism. As the size of the data increases, both our DGKT framework and DGrKT consistently maintains its advantage.
Moreover, we compare our DGrKT with existing KT models with different data size, as shown in Fig. 7. We find that as data size decreasing, other KT models exhibit a significant drop in performance. However, the performance of DGrKT declines at a much slower rate, indicating that it is highly adaptable to limited data size.
Figure 6: The AUC result for three existing KT models (DKT [28], AKT [15], SANIT [10]) on ASSISTment 2009 with different data sizes compared to DGrKT.
Figure 7: The AUC result for different $\lambda$ with different data sizes. Each colored line represents the AUC as $\lambda$ changes under a specific data size setting.
4.3 Ablation Studies.
Table 3: The AUC results for different varints of DGrKT on five target domain datasets.
Models batch num ASSIST09 ASSIST15 ASSIST17 ALGEBRA Junyi
DGrKT -w/o CFL 1 0.699 0.670 0.624 0.701 0.691
2 0.705 0.689 0.649 0.756 0.704
4 0.716 0.707 0.667 0.764 0.709
8 0.726 0.719 0.675 0.772 0.713
DGrKT -w/o CR 1 0.711 0.697 0.646 0.750 0.700
2 0.717 0.709 0.663 0.762 0.706
4 0.725 0.718 0.672 0.767 0.712
8 0.728 0.723 0.681 0.774 0.714
DGrKT -w/o TA 1 0.709 0.695 0.645 0.752 0.697
2 0.718 0.712 0.662 0.763 0.707
4 0.724 0.717 0.672 0.767 0.711
8 0.729 0.723 0.680 0.780 0.713
DGrKT 1 0.729 0.717 0.677 0.775 0.724
2 0.737 0.729 0.684 0.781 0.727
4 0.740 0.732 0.689 0.784 0.728
8 0.743 0.732 0.696 0.795 0.730
Concept Aggregation. We conducted ablation experiments to investigate the impact of our proposed concept aggregation methods, which include concept feature learning (CFL), concept prototypes refinement (CR), and target concept embedding adaptation (TA). Specifically, we compare our DGrKT model with three variants: DGrKT-w/o CFL, DGrKT-w/o CR, and DGrKT-w/o TA. The details of these variants are presented below:
  • 1.
    DGrKT-w/o CFL: This variant omits concept feature learning. The model is directly trained on the target domain without pretraining on the source domain.
  • 2.
    DGrKT-w/o CR: This variant omits concept prototypes refinement. After concept feature learning, the model is directly fine-tuned on the target domain without concept prototypes refinement.
  • 3.
    DGrKT-w/o TA: This variant omits target concept embedding adaptation. The model is fine-tuned on the target domain using randomly initialized embeddings, and question embeddings are calculated by Eqn. (24).
From Table 3, we can observe that DGrKT without Concept Feature Learning (CFL) shows a significant decrease in AUC. This indicates that CFL is a crucial procedure, as it derives meta-knowledge from several source domains. Additionally, concept aggregation proves to be beneficial for learning a robust concept embedding for DGrKT, as well as for adapting the target concept embeddings effectively.
Table 4: The AUC results for different varints of DG-DKT on five target domain datasets.
Models batch num ASSIST09 ASSIST15 ASSIST17 ALGEBRA Junyi
- 1 0.712 0.689 0.657 0.715 0.634
2 0.717 0.700 0.677 0.743 0.653
4 0.726 0.704 0.680 0.764 0.681
8 0.731 0.706 0.690 0.765 0.682
BN 1 0.718 0.691 0.671 0.748 0.668
2 0.722 0.702 0.679 0.759 0.669
4 0.730 0.710 0.691 0.762 0.680
8 0.729 0.715 0.697 0.775 0.675
LN 1 0.713 0.688 0.651 0.746 0.749
2 0.725 0.704 0.671 0.752 0.760
4 0.727 0.713 0.679 0.760 0.764
8 0.730 0.719 0.695 0.772 0.776
SeqIN 1 0.721 0.694 0.679 0.749 0.678
2 0.727 0.705 0.689 0.760 0.684
4 0.732 0.710 0.695 0.764 0.689
8 0.734 0.716 0.703 0.776 0.691
Sequence Instance Normalization. Moreover, we conduct ablation experiment on Sequence Instance Normalization (SeqIN). Thus, we compare DG-DKT with one variant without any normalization method and two variants with BN and LN respectively.
We omit instance normalization because its calculation inevitably includes all timesteps, leading to information leakage in knowledge tracing. The results, shown in Table 4, demonstrate the advantage of SeqIN. Specifically, DG-DKT without SeqIN (DG-DKT -w/o SeqIN) shows a significant decrease in AUC, indicating that the normalization method is crucial in such cross-domain tasks and Sequence Instance Normalization is a better normalization module for multi-source domain adaptation knowledge tracing.
4.4 Visualization
@@
[…]
@@
Figure 8: The $t$-SNE visualization of feature embedding of student interactions by using different normalization methods. Different colors represent different source domains.
Figure 9: A visualization of a student’s mastery across five concept prototypes, based on an interaction randomly selected from ASSIST09. Different colors represent distinct concept prototypes, with hollow circles indicating incorrect responses and filled circles indicating correct responses for the corresponding concept prototype.
[…]
Figure 10: Visualization of attention scores from DGrKT and DGrKT-w/o RA. As shown, the attention scores in DGrKT concentrate more effectively on relevant timesteps, while in DGrKT-w/o RA, the attention is spread more uniformly across all timesteps, lacking focus on the crucial related timesteps.
Visualization of student’s mastery. We also pivot towards gauging a student’s mastery of five concept clusters via visualization, which is calculated by his predicted performance. The interaction sequence is derived from ASSIST09, consisting of the initial 50 interactions, as is shown in Fig. 9. A green rectangle implies that the student correctly answered the corresponding concept cluster while red rectangle means that the student failed. We find that our DGrKT is capable of providing a coarse-grained but effective assessment despite the data scarcity issue.
Visualization of attention score. Here, we demonstrate the effectiveness of relation-aware attention using a heatmap visualization. As shown in Fig. 10, we present the attention scores for four attention heads: the left side displays scores from DGrKT without relation-aware attention (DGrKT -w/o RA), while the right side shows scores from DGrKT with relation-aware attention. We observe that with the help of relation-aware attention, DGrKT effectively distinguishes the relationships between different timesteps, focusing on those with closer connections. In contrast, the attention scores in DGrKT -w/o RA are smoother, indicating a reliance primarily on the sequential order.
Cluster analysis. We conducted an analysis on the aggregated concepts from ASSIST09, randomly selecting 8 concepts from each cluster. As shown in Table 5, it is evident that the first cluster primarily consists of fundamental and basic concepts, while the second cluster contains many advanced algebraic concepts. The third, fourth, and last clusters respectively include geometry concepts, advanced data representation, and statistical concepts. Although the concepts are not strictly categorized into five clusters, those with similar patterns in student interactions have been grouped together. These results clearly demonstrate the effectiveness of the concept aggregation process.
Table 5: Visualization of concept aggregation results.
@@
4 Venn Diagram, Counting Methods, Congruence, Proportion, Prime Number, Algebraic Simplification, Equation Solving More Than Two Steps, Write Linear Equation from Situation